### v2

@@ 4.2 Deriving the strong second-order sufficient condition
On the other hand, if there exists a vector $\tilde{\bm{\xi}}\in-{\sf int}{\mathcal{Q}}_{j_{\iota}}$ such that $\langle\tilde{\bm{\xi}},M\tilde{\bm{\xi}}\rangle\leq 0$, we obtain from (4.46) and Lemma 2.1 that there exists a nonzero vector ${\boldsymbol{v}}\in-{\sf int}{\mathcal{Q}}_{j_{\iota}}$ such that $M{\boldsymbol{v}}\in{\mathcal{Q}}_{j_{\iota}}$. Then, one has $\langle{\boldsymbol{v}},M{\boldsymbol{v}}\rangle\leq 0$, i.e., $\langle(M{\boldsymbol{v}}),M^{-1}(M{\boldsymbol{v}})\rangle\leq 0$. This contradicts either (4.43) or (4.44) since ${\boldsymbol{0}}\neq M{\boldsymbol{v}}\in{\mathcal{Q}}_{j_{\iota}}$. Thus (4.45) holds. Combining (4.43), (4.44), and (4.45) we know from Lemma 2.2 that (4.40) does not hold. This completes the proof. ∎
@@
Let ${\boldsymbol{x}}^{*}$ be a locally optimal solution to the nonlinear SOCP (1.1) with ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ being an associated multiplier. Let the index set $\{1,\ldots,J\}$ be classified as in (4.1). Assume that the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ given in (1.5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$. Then the strong second-order sufficient condition (4.6) holds for (1.1) at ${\boldsymbol{x}}^{*}$.
Proof.
Since the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, we know that the mapping ${\mathbb{S}}_{\rm GE}$ in (1.8) also has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$. Then, by [10, Corollary 25] the constraint nondegeneracy condition (1.7) holds at ${\boldsymbol{x}}^{*}$ and the second-order sufficient condition holds (4.5) at ${\boldsymbol{x}}^{*}$. Therefore, Assumption 4.1 holds.
Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix defined in (4.12), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. From Corollary 4.1 we know that (4.17) holds with ${\mathcal{G}}_{j}$ ($j\in J_{4,5,6}$) being defined by (4.18). Furthermore, one knows that (4.20) holds by Corollary 4.2. Then, based on (4.24) and (4.25), one has (4.30) holds for the matrices $R_{i}$, $P_{i}$ and $T_{i}$ defined in (4.26), (4.28) and (4.29) by Proposition 4.2.
From Lemma 4.3 we know that $\langle{\bm{\nu}},H^{\top}\Theta H{\bm{\nu}}\rangle>0,\,\forall{\bm{\nu}}\in{\mathcal{V}}\setminus\{{\boldsymbol{0}}\}$, where ${\mathcal{V}}$ is defined in (4.22). Therefore, one has
@@
${\bm{\nu}}\in{\sf rge}P_{|J_{4,5,6}|}.$ (4.48)
On the other hand, if (4.48) holds, then it is easy to see from (4.27) that ${\bm{\nu}}\in{\sf lin}{\mathcal{V}}$ holds. Therefore, ${\bm{\nu}}\in{\sf lin}{\mathcal{V}}$ if and only if (4.48) holds. Consequently, by (4.47) one has $\langle{\bm{\nu}},H^{\top}\Theta H{\bm{\nu}}\rangle>0$ for all ${\boldsymbol{0}}\neq{\bm{\nu}}\in{\sf rge}P_{|J_{4,5,6}|}$. This is equivalent, by (4.28) and (4.29), to the condition that
$T_{i}=P_{i}^{\top}H^{\top}\Theta HP_{i}=R_{i}^{\top}T_{i-1}R_{i}\succ O\quad\mbox{with}\quad i=|J_{4,5,6}|.$ (4.49)
Recall that each $R_{i}$ defined in (4.26) has full column rank and each $P_{i}$ in (4.28) also has full column rank. From Corollary 4.2 we know that each ${\mathcal{G}}_{j_{i}}$ is surjective. Then by (4.27) we know that
$({\sf rge}\,R_{i})^{\circ}={\sf rge}[{\mathcal{G}}_{j_{i}}P_{i-1}]^{\top},\quad\forall\,i=1,\ldots,|J_{4,5,6}|.$ (4.50)
Recall from (4.30) of Proposition 4.2 that
$[{\mathcal{G}}_{{j_{i}}}P_{i-1}]T_{i-1}^{-1}[{\mathcal{G}}_{j_{i}}P_{i-1}]^{\top}\succ O\quad\mbox{with}\quad i=|J_{4,5,6}|.$ (4.51)
Based on (4.49), (4.50) and (4.51), one can apply Lemma 2.2 to get $T_{|J_{4,5,6}|-1}\succ O$. Furthermore, for any index $i\in\{2,\ldots,|J_{4,5,6}|-1\}$, if one has
$T_{i}=R_{i}^{\top}T_{i-1}R_{i}\succ O,$ (4.52)
one can get from (4.30) of Proposition 4.2 that $[{\mathcal{G}}_{j_{i}}P_{i-1}]T_{i-1}^{-1}[{\mathcal{G}}_{j_{i}}P_{i-1}]^{\top}\succ O$, which, together with (4.50) and (4.52), implies $T_{i-1}\succ O$ by Lemma 2.2. Therefore, one can get $T_{0}=H^{\top}\Theta H\succ O$ by induction. Note that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$ and $\Theta\in{\mathds{R}}^{n\times n}$ is the matrix defined in (4.12). Then from Lemma 4.3 we know that the strong second-order sufficient condition (4.6) holds. ∎
@@ 4.3 The equivalence theorem
Let ${\boldsymbol{x}}^{*}$ be a locally optimal solution to the nonlinear SOCP (1.1) and ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ be a corresponding multiplier. The following two statements are equivalent:
@@
    the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$;
@@
    the solution $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is strongly regular to the KKT system (1.6).
Proof.
Note that (1) is an immediate consequence of (2). On the other hand, if (1) holds, we know from [10, Corollary 25] that ${\boldsymbol{x}}^{*}$ is nondegenerate and the second-order sufficient condition (4.5) holds at ${\boldsymbol{x}}^{*}$. Consequently, it follows from Proposition 4.3 that the strong second-order sufficient condition (4.6) holds. Thus, by using [6, Theorem 30] we know that (2) holds. This completes the proof. ∎
Theorem 4.1 establishes the equivalence between the Aubin property and the strong regularity for the nonlinear SOCP (1.1) at local optimal solutions without requiring the strict complementarity. Moreover, Theorem 4.1 implies its counterpart for conventional nonlinear programming (((([12, Theorems 1, 4 &\&& 5])))) as a special case as one can take $r_{j}=0$ for all $j=1,\dots,J$.
According to Theorem 4.1, ${{\mathbb{S}}}_{\rm KKT}$ in (1.5) having the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is also equivalent to many other conditions, such as (a) ${\boldsymbol{x}}^{*}$ is nondegenerate and the uniform second order growth condition ([5, Definition 5.16]) holds at ${\boldsymbol{x}}^{*}$ (by [6, Theorem 30]); (b) ${\boldsymbol{x}}^{*}$ is a nondegenerate and fully stable ([24, Definition 4.1]) locally optimal solution (by [24, Theorem 4.8]); and (c) Clarke’s generalized Jacobian of a KKT system is nonsingular (by [37, Theorem 3.1]). One may refer to the relevant references for details.

### v3

@@ 4.2 Deriving the strong second-order sufficient condition
On the other hand, if there exists a vector $\tilde{\bm{\xi}}\in-{\sf int}{\mathcal{Q}}_{j_{\iota}}$ such that $\langle\tilde{\bm{\xi}},M\tilde{\bm{\xi}}\rangle\leq 0$, we obtain from (72) and Lemma 2.1 that there exists a nonzero vector ${\boldsymbol{v}}\in-{\sf int}{\mathcal{Q}}_{j_{\iota}}$ such that $M{\boldsymbol{v}}\in{\mathcal{Q}}_{j_{\iota}}$. Then, one has $\langle{\boldsymbol{v}},M{\boldsymbol{v}}\rangle\leq 0$, i.e., $\langle(M{\boldsymbol{v}}),M^{-1}(M{\boldsymbol{v}})\rangle\leq 0$. This contradicts either (69) or (70) since ${\boldsymbol{0}}\neq M{\boldsymbol{v}}\in{\mathcal{Q}}_{j_{\iota}}$. Thus (71) holds. Combining (69), (70), and (71) we know from Lemma 2.3 that (66) does not hold. This completes the proof.
@@
Let ${\boldsymbol{x}}^{*}$ be a locally optimal solution to the nonlinear SOCP (1) with ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ being an associated multiplier. Let the index set $\{1,\ldots,J\}$ be classified as in (27). Assume that the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ given in (5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$. Then the strong second-order sufficient condition (32) holds for (1) at ${\boldsymbol{x}}^{*}$.
Proof 4.18.
Since the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$, we know that the mapping ${\mathbb{S}}_{\rm GE}$ in (8) also has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$. Then, by [10, Corollary 25] the constraint nondegeneracy condition (7) holds at ${\boldsymbol{x}}^{*}$ and the second-order sufficient condition holds (31) at ${\boldsymbol{x}}^{*}$. Therefore, Assumption 4.1 holds.
Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix defined in (38), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. From Corollary 4.9 we know that (43) holds with ${\mathcal{G}}_{j}$ ($j\in J_{4,5,6}$) being defined by (44). Furthermore, one knows that (46) holds by Corollary 4.11. Then, based on (50) and (51), one has (56) holds for the matrices $R_{i}$, $P_{i}$ and $T_{i}$ defined in (52), (54) and (55) by Proposition 4.15.
From Lemma 4.13 we know that $\langle{\bm{\nu}},H^{\top}\Theta H{\bm{\nu}}\rangle>0,\,\forall{\bm{\nu}}\in{\mathcal{V}}\setminus\{{\boldsymbol{0}}\}$, where ${\mathcal{V}}$ is defined in (48). Therefore, one has
@@
(74) ${\bm{\nu}}\in{\sf rge}P_{|J_{4,5,6}|}.$
On the other hand, if (74) holds, then it is easy to see from (53) that ${\bm{\nu}}\in{\sf lin}{\mathcal{V}}$ holds. Therefore, ${\bm{\nu}}\in{\sf lin}{\mathcal{V}}$ if and only if (74) holds. Consequently, by (73) one has $\langle{\bm{\nu}},H^{\top}\Theta H{\bm{\nu}}\rangle>0$ for all ${\boldsymbol{0}}\neq{\bm{\nu}}\in{\sf rge}P_{|J_{4,5,6}|}$. This is equivalent, by (54) and (55), to the condition that
(75) $T_{i}=P_{i}^{\top}H^{\top}\Theta HP_{i}=R_{i}^{\top}T_{i-1}R_{i}\succ O\quad\mbox{with}\quad i=|J_{4,5,6}|.$
Recall that each $R_{i}$ defined in (52) has full column rank and each $P_{i}$ in (54) also has full column rank. From Corollary 4.11 we know that each ${\mathcal{G}}_{j_{i}}$ is surjective. Then by (53) we know that
(76) $({\sf rge}\,R_{i})^{\circ}={\sf rge}[{\mathcal{G}}_{j_{i}}P_{i-1}]^{\top},\quad\forall\,i=1,\ldots,|J_{4,5,6}|.$
Recall from (56) of Proposition 4.15 that
(77) $[{\mathcal{G}}_{{j_{i}}}P_{i-1}]T_{i-1}^{-1}[{\mathcal{G}}_{j_{i}}P_{i-1}]^{\top}\succ O\quad\mbox{with}\quad i=|J_{4,5,6}|.$
Based on (75), (76) and (77), one can apply Lemma 2.3 to get $T_{|J_{4,5,6}|-1}\succ O$. Furthermore, for any index $i\in\{2,\ldots,|J_{4,5,6}|-1\}$, if one has
(78) $T_{i}=R_{i}^{\top}T_{i-1}R_{i}\succ O,$
one can get from (56) of Proposition 4.15 that $[{\mathcal{G}}_{j_{i}}P_{i-1}]T_{i-1}^{-1}[{\mathcal{G}}_{j_{i}}P_{i-1}]^{\top}\succ O$, which, together with (76) and (78), implies $T_{i-1}\succ O$ by Lemma 2.3. Therefore, one can get $T_{0}=H^{\top}\Theta H\succ O$ by induction. Note that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$ and $\Theta\in{\mathds{R}}^{n\times n}$ is the matrix defined in (38). Then from Lemma 4.13 we know that the strong second-order sufficient condition (32) holds.
@@ 4.3 The equivalence theorem
Let ${\boldsymbol{x}}^{*}$ be a locally optimal solution to the nonlinear SOCP (1) and ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ be a corresponding multiplier. The following two statements are equivalent:
@@
    the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ in (5) has the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$;
@@
    the solution $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is strongly regular to the KKT system (6).
Proof 4.20.
Note that (1) is an immediate consequence of (2). On the other hand, if (1) holds, we know from [10, Corollary 25] that ${\boldsymbol{x}}^{*}$ is nondegenerate and the second-order sufficient condition (31) holds at ${\boldsymbol{x}}^{*}$. Consequently, it follows from Proposition 4.17 that the strong second-order sufficient condition (32) holds. Thus, by using [6, Theorem 30] we know that (2) holds. This completes the proof.
Theorem 4.19 establishes the equivalence between the Aubin property and the strong regularity for the nonlinear SOCP (1) at local optimal solutions without requiring the strict complementarity. Moreover, Theorem 4.19 implies its counterpart for conventional nonlinear programming (((([12, Theorems 1, 4 &\&& 5])))) as a special case as one can take $r_{j}=0$ for all $j=1,\dots,J$.
According to Theorem 4.19, ${{\mathbb{S}}}_{\rm KKT}$ in (5) having the Aubin property at $({\boldsymbol{0}},{\boldsymbol{0}})$ for $({\boldsymbol{x}}^{*},{\boldsymbol{y}}^{*})$ is also equivalent to many other conditions, such as (a) ${\boldsymbol{x}}^{*}$ is nondegenerate and the uniform second order growth condition ([5, Definition 5.16]) holds at ${\boldsymbol{x}}^{*}$ (by [6, Theorem 30]); (b) ${\boldsymbol{x}}^{*}$ is a nondegenerate and fully stable ([24, Definition 4.1]) locally optimal solution (by [24, Theorem 4.8]); and (c) Clarke’s generalized Jacobian of a KKT system is nonsingular (by [37, Theorem 3.1]). One may refer to the relevant references for details.
@@ 5 Conclusions
Acknowledgments
The authors would like to thank the two anonymous referees and the associate editor for their many valuable comments and constructive suggestions including the references mentioned in their reports, which have substantially helped improve the quality and presentation of this paper.